f(x) = x + 7
g(x) = sqrt (x)

f + g
f - g
fg
f/g

Answer by edjones(8007)

f + g = x+7+sqrt(x)
f - g = x+7-sqrt(x)
fg = sqrt(x)(x+7) = x*sqrt(x)+7sqrt(x)
f/g = (x+7)/sqrt(x) = (x+7)/sqrt(x) * sqrt(x)/sqrt(x) = sqrt(x)(x+7)/x
.
Ed